OpenStudy (anonymous):

Find the derivative of f(x) = -6/x at x = 12.

OpenStudy (anonymous):

whats the derivative?

OpenStudy (anonymous):

the derivative is -6/x^2

OpenStudy (anonymous):

at x=12 i will be -6/144=-1/24

OpenStudy (anonymous):

-6*x^(-1), formula for variables with a power x^n= (x^(n-1))/(n-1)

OpenStudy (anonymous):

*n-1 not divided

OpenStudy (anonymous):

6/x^2

OpenStudy (anonymous):

is derivative

OpenStudy (anonymous):

i think is positive ikarus

OpenStudy (anonymous):

its*

OpenStudy (anonymous):

These are the choices 24 1/2 1/24 2

OpenStudy (anonymous):

third

OpenStudy (anonymous):

ok 1/24

OpenStudy (anonymous):

yeah it's positive @Lost_in_Translation

OpenStudy (anonymous):

sorry i missed that

OpenStudy (anonymous):

not a problem signs are evil